OpenStudy (anonymous):

What is the sum of 5√7 and 3√28 ?

OpenStudy (anonymous):

\[28=4\times 7\] so \[\sqrt{28}=\sqrt{4}\times \sqrt{7}=2\sqrt{7}\]

OpenStudy (anonymous):

and \[5\sqrt{7}+2\sqrt{7}=7\sqrt{7}\]